### v36

# Weighing the Axion with Muon Haloscopy

Abstract
Recent measurements of muon spin precession confirm a long-standing tension with the Standard Model of particle physics. We argue that axions from the local dark matter halo of the galaxy are responsible for the tension. The argument yields a percent level prediction for the mass of the axion provided that dark matter is made of axions. An analysis of charge asymmetry in kaon decays suggests that at least in the local halo dark matter is made of axions and that axions from the local halo are responsible for the observed violation by these reactions of the combined charge conjugation and spatial inversion symmetry operation. A tabletop experiment to directly detect dark matter in the form of axions with the predicted mass is proposed.
Introduction—The spin of the muon precesses in a magnetic field $B$ at frequency $\omega_{a}=a_{\mu}(e/m_{\mu})B$, where, $a_{\mu}\approx\alpha/(2\pi)\approx 1.16\times 10^{-3}$ is the electromagnetic leading order contribution to spin precession, and $e/m_{\mu}=2\pi\times 135~{}{\rm MHz/T}$ is the charge to mass ratio of the muon [1, 2, 3, 4, 5]. In this Letter, we show that axions in the local dark matter halo of the galaxy shift $a_{\mu}$ from the standard model value. The demonstration rests on a novel physical picture for the nature of dark matter which leads to a percent level prediction for the mass of the axion.
Nature of Dark Matter—We start by expressing the axion $A_{M}$ in terms of quarks and leptons $M^{P}_{H}$ with helicity $H=L,R$ and charge-parity check $P=+,-$ [6]:
$A_{M}$ $=$ $M_{L}^{-}\overline{M}_{R}^{-}-M_{R}^{-}\overline{M}_{L}^{-}-M_{L}^{+}\overline{M}_{R}^{+}+M_{R}^{+}\overline{M}_{L}^{+}$ (1)
Here, $M$ runs over the twelve known “flavors” of quarks and leptons while the charge-parity check $P$ is $+$ for quarks and leptons that do not feel the weak nuclear force and -- for those that do. In the standard model, we find $M_{L}^{-}$ and $M_{R}^{+}$, but neither $M_{L}^{+}$ nor $M_{R}^{-}$, where, $L$ is for left-handed and $R$ is for right-handed helicity [7].
Next, we suppose that axions $A_{M}$ of each kind and photons $\gamma_{H}$ of each helicity formed in equal numbers in the early universe prior to baryogenesis. This fixes the ratio $n_{A}/n_{\gamma}=|M|/|H|=6$ of axions to photons and gives the axion mass $m_{A}$ ($\hbar=c=1$) [8]:
$m_{A}=2.70~{}kT_{\gamma}\left(\frac{n_{A}}{n_{\gamma}}\right)^{-1}\frac{\Omega_{A}h^{2}}{\Omega_{\gamma}h^{2}}=(0.508\pm 0.004)~{}{\rm eV},$ (2)
where, $T_{\gamma}=(2.7255\pm 0.0006)~{}{\rm K}$ is the present photon temperature [9], $\Omega_{\gamma}h^{2}=2.473\times 10^{-5}$ is the photon energy density parameter and $\Omega_{A}h^{2}=0.11882\pm 0.00086$ is that of dark matter [10].
Local Dark Matter Halo—The energy density of axions in the local dark matter halo of the galaxy $\rho_{A}$ can be estimated using charge asymmetry of kaon decays [11]:
$A_{L}(e)$ $=$ $-2\sqrt{2}\xi_{A\pi}p_{K}\tilde{\phi}_{A}(q_{A})$ (3)
$=$ $4\pi C_{A\pi}\frac{f_{\pi}\sqrt{\rho_{A}hc}}{m_{A}f_{A}}L_{K},$
where, the axion amplitude $\tilde{\phi}_{A}(q_{A})$ at momentum $q_{A}$ changes sign under the combination of charge conjugation and space inversion probed by $A_{L}(e)$ to create the observed charge asymmetry [6]. The mixing angle $\xi_{A\pi}=-C_{A\pi}f_{\pi}/f_{A}$ between the axion $A$ and the neutral pion $\pi^{0}$ can be expressed in terms of the pion decay constant $f_{\pi}=(92.32\pm 0.09)~{}{\rm MeV}$ [12], the axion decay constant $f_{A}$ [13], and the light quark masses using $C_{A\pi}=(m_{d}-m_{u})/(2m_{u}+2m_{d})=(0.173\pm 0.008)$ [14]. Here, the kaon beam has momentum $p_{K}$ and decays in-flight within a detector of length $L_{K}$.
Figure 1: Halo axion $A$ collides with virtual photon dressing vertex of positive muon $\mu^{+}$ and magnetic field $B$.
In contrast to the local axion energy density, the axion mass $m_{A}$ and decay constant $f_{A}$ only enter the kaon charge asymmetry through the product $m_{A}f_{A}=\sqrt{\chi_{\rm QCD}}=(5.69\pm 0.05)\times 10^{-3}~{}{\rm GeV^{2}}$ which is fixed by the topological susceptibility $\chi_{\rm QCD}$ of the quantum chromodynamic vacuum [15]. Using the kaon parameters $L_{K}=500~{}{\rm cm}$ and $A_{L}(e)=(3.32\pm 0.08)\times 10^{-3}$ from the current best charge asymmetry measurement [16, 17], the local axion energy density $\rho_{A}=(0.28\pm 0.03)~{}{\rm GeV/cm^{3}}$ that we find from Eq. (3) agrees with the best available estimate $\rho_{\rm DM}=(0.30\pm 0.03)~{}{\rm GeV/cm^{3}}$ of the energy density of the local dark matter halo [18]. This agreement suggests that axions saturate the local dark matter halo and that dark matter from the local halo is the source of the charge asymmetry observed in kaon decays.
Muon Haloscopy—Figure 1 shows how axions in the local dark matter halo shift muon spin precession [3, 4, 5]. Acting as a background field $\tilde{\phi}_{A}(q_{A})$ with on-shell four-momentum $q_{A}^{2}=m_{A}^{2}\ll m_{\mu}^{2}$, the axion $A$ couples to the virtual photon dressing the vertex of the positive muon $\mu^{+}$ and the magnetic field $B$. The axion-photon coupling of strength $g_{A\gamma\gamma}=(0.68\pm 0.02)\times 10^{-10}~{}{\rm GeV^{-1}}$ [19] changes the photon’s Green’s function [20, 21]
$\Delta D_{+}(x,x^{\prime})=m_{A}^{2}g_{A\gamma\gamma}\tilde{\phi}(q_{A})D_{+}(x,x^{\prime}),$ (4)
and this change in the photon propagator produces the shift of the spin precession frequency [22]:
$\frac{\Delta a_{\mu}}{a_{\mu}}$ $=$ $m_{A}^{2}g_{A\gamma\gamma}\tilde{\phi}_{A}(q_{A})$ (5)
$=$ ±$\pm m_{A}g_{A\gamma\gamma}\sqrt{\rho_{A}N_{\mu}V_{\mu}T_{\mu}}$
where, the axion field strength $|\tilde{\phi}_{A}(q_{A})|^{2}\approx\phi_{A}^{2}N_{\mu}V_{\mu}T_{\mu}$ is fixed by the axion energy density $\rho_{A}=\phi_{A}^{2}m_{A}^{2}$, the volume $V_{\mu}$ of the muon beam, the lifetime $T_{\mu}$ of the muon in the axion rest-frame, and the average number of muons $N_{\mu}$ during spin precession.
The sign of the axion-driven shift $\Delta a_{\mu}$ of the muon spin precession frequency switches along with the axion field $\tilde{\phi}_{A}(q_{A})$ under spatial inversion $P$, under time-reversal $T$, and under the combinations $CP$ and $CT$ with charge conjugation $C$ [23]. The size of the shift scales with muon beam parameters [24]
$\frac{\Delta a_{\mu}}{a_{\mu}}$ $=$ $(165\pm 10)\times 10^{-11}\left(\frac{N_{\mu}}{1930}\right)^{1/2}\left(\frac{\sigma_{x}}{1.78~{}{\rm cm}}\right)^{1/2}$
$\times$ $\left(\frac{\sigma_{y}}{1.25~{}{\rm cm}}\right)^{1/2}\left(\frac{R}{711~{}{\rm cm}}\right)^{1/2}\left(\frac{T_{\mu}}{64.4~{}{\rm\mu sec}}\right)^{1/2}$
where, $\sigma_{x}$ is the width of the muon beam, $\sigma_{y}$ is its height, and $R$ is the radius of the storage ring. This has roughly the right size to resolve the $3.4~{}\sigma$ tension between the spin precession experiments (Exp) [2, 3, 4] and standard model calculations (SM) [25, 26, 27, 28, 29, 30]
$a_{\mu}({\rm Exp})-a_{\mu}({\rm SM})=(140\pm 41)\times 10^{-11},$ (7)
where, we have used the recently emerged consensus value $a_{\mu}^{\rm HLO}({\rm SM})=(704.1\pm 3.0)\times 10^{-10}$ [31] for the leading order hadronic contribution to the muon spin precession frequency within the Standard Model [32]. Tabletop Dark Matter Detector— Axions from the local dark matter halo of the galaxy can be seen in the mid-infrared electromagnetic response of polar materials such as silicon nitride ${\rm Si_{3}N_{4}}$ [33]. The incoming mid-infrared light couples to the axion and the coupling creates far-infrared electromagnetic radiation at the resonance frequency $\omega_{\theta}$ of the transverse optical phonon that dominates the far-infrared response of the polar material. The mid-infrared response hits the axion resonance at frequency $\omega_{A\theta}=\omega_{A}-\omega_{\theta}$ which is detuned by the frequency of the transverse optical phonon $\omega_{\theta}$ from the frequency $\omega_{A}=m_{A}c^{2}/\hbar=2\pi\times(122.7\pm 1.0)~{}{\rm THz}=2\pi c\times(4097\pm 32)~{}{\rm cm^{-1}}$ set by the axion rest-mass.
Measuring the mid-infrared frequency $\omega_{A\theta}=\omega_{A}-\omega_{\theta}$ that resonantly excites the polar material provides a table-top determination of the mass of the axion. The size of the dip in mid-infrared transmission through a sample of the polar material on resonance at $\omega_{A\theta}$ is set by the infrared absorption coefficient [34]:
$\alpha(\omega_{A\theta})$ $=$ $\alpha(\omega_{\theta})\frac{\Gamma_{A\theta}}{\Gamma_{\theta}}\frac{\omega_{\theta}}{\omega_{A\theta}},$ (8)
where $\alpha(\omega_{\theta})$ is the absorption coefficient of the phonon, $\Gamma_{A\theta}\approx 2\pi c/t$ gives the width of the axion resonance in terms of the thickness of the sample $t$, and $\Gamma_{\theta}$ is the width of the phonon resonance. Using typical parameters for a polar material such as silicon nitride ${\rm Si_{3}N_{4}}$, we find the mid-infrared absorption coefficient at $\omega_{A\theta}$ from the axion resonance is large enough to be readily observable with samples that show a dip in the far-infrared at frequency $\omega_{\theta}$ from the transverse optical phonon [33]
$\alpha(\omega_{A\theta})$ $=$ $4.0\times\alpha(\omega_{\theta})\left(\frac{t}{5~{}{\rm\mu m}}\right)^{-1}\left(\frac{\Gamma_{\theta}/2\pi c}{10~{}{\rm cm^{-1}}}\right)^{-1}$ (9)
$\times$ $\left(\frac{\omega_{\theta}/2\pi c}{80~{}{\rm cm^{-1}}}\right)\left(\frac{\omega_{A\theta}/2\pi c}{4016~{}{\rm cm^{-1}}}\right)^{-1}.$
The mid-infrared absorption in polar materials such as silicon nitride ${\rm Si_{3}N_{4}}$ at frequency $\omega_{A\theta}=\omega_{A}-\omega_{\theta}$ due to axions from the local dark matter halo of the galaxy can thus be seen with readily available samples and infrared spectrometers [33]
Conclusion— We conclude that the tensions between measurements of muon spin precession and standard model calculations are due to axions from the local dark matter halo of the galaxy. The resolution of the muon tension rests on a novel physical picture for the nature of dark matter. The picture leads to percent level predictions for the mass of the axion and for its coupling to electromagnetic fields which can be tested using a tabletop experiment with readily available materials and devices.
Acknowledgements— This work was performed in part at the Kavli Institute for Theoretical Physics, which is supported by the National Science Foundation grant No. PHY-1748958, at the Aspen Center for Physics, which is supported by National Science Foundation grant No. PHY-2210452, at the Institute for Nuclear Theory, which is supported by the Department of Energy under grant No. DE-FG02- 00ER41132, and at the Mainz Institute of Theoretical Physics, which is part of the Cluster of Excellence PRISMA+ Project ID 39083149.