### v37

# Weighing the Axion with Muon Haloscopy

Abstract
Recent measurements of muon spin precession confirm a long-standing tension with the standard model of particle physics. We argue that axions form the dark matter in the universe and that axions from the local dark matter halo of the galaxy are responsible for the tension. The argument gives a percent level prediction for the mass of the axion that can be tested by measuring the mid-infrared electrodynamic response of a wide range of polar materials. A broad resonance in the scattered visible light intensity from quartz crystal at room-temperature and ambient pressure is reported that peaks at a frequency shift of the light which is consistent with the creation of axions with the predicted mass.
Introduction—The spin of the muon precesses in a magnetic field $B$ at frequency $\omega_{a}=a_{\mu}(e/m_{\mu})B$, where, $a_{\mu}\approx\alpha/(2\pi)\approx 1.16\times 10^{-3}$ is the electromagnetic leading order contribution to spin precession, and $e/m_{\mu}=2\pi\times 135~{}{\rm MHz/T}$ is the charge to mass ratio of the muon [1, 2, 3, 4, 5]. In this Letter, we show that axions in the local dark matter halo of the galaxy shift $a_{\mu}$ from the standard model value. The demonstration rests on a novel physical picture for the nature of dark matter which leads to a percent level prediction for the mass of the axion.
Nature of Dark Matter—We start by expressing the axion $A_{M}$ in terms of quarks and leptons $M^{P}_{H}$ with helicity $H=L,R$ and charge-parity check $P=+,-$ [6]:
$A_{M}$ $=$ $M_{L}^{-}\overline{M}_{R}^{-}-M_{R}^{+}\overline{M}_{L}^{+}-M_{L}^{+}\overline{M}_{R}^{+}+M_{R}^{-}\overline{M}_{L}^{-}$ (1)
Here, $M$ runs over the twelve known “flavors” of quarks and leptons while the charge-parity check $P$ is $+$ for quarks and leptons that do not feel the weak nuclear force and -- for those that do. In the standard model, we find $M_{L}^{-}$ and $M_{R}^{+}$, but neither $M_{L}^{+}$ nor $M_{R}^{-}$, where, $L$ is for left-handed and $R$ is for right-handed helicity [7].
Next, we suppose that axions $A_{M}$ of each kind and photons $\gamma_{H}$ of each helicity formed in equal numbers in the early universe prior to baryogenesis. This fixes the ratio $n_{A}/n_{\gamma}=|M|/|H|=6$ of axions to photons and gives the axion rest-mass energy [8]:
$m_{A}c^{2}=2.70~{}kT_{\gamma}\left(\frac{n_{A}}{n_{\gamma}}\right)^{-1}\frac{\Omega_{A}h^{2}}{\Omega_{\gamma}h^{2}}=(0.508\pm 0.004)~{}{\rm eV},$ (2)
where, $T_{\gamma}=(2.7255\pm 0.0006)~{}{\rm K}$ is the present photon temperature [9], $\Omega_{\gamma}h^{2}=(2.473\pm 0.002)\times 10^{-5}$ is the photon energy density parameter and $\Omega_{A}h^{2}=0.11882\pm 0.00086$ is that of dark matter [10].
Figure 1: Halo axion $A$ collides with virtual photon dressing vertex of positive muon $\mu^{+}$ and magnetic field $B$.
Muon Haloscopy—Figure 1 shows how axions in the local dark matter halo shift muon spin precession [3, 4, 5]. Acting as a background field $\tilde{\phi}_{A}(q_{A})$ with on-shell four-momentum $q_{A}^{2}=(m_{A}c)^{2}\ll(m_{\mu}c)^{2}$, the axion $A$ couples to the virtual photon dressing the vertex of the positive muon $\mu^{+}$ and the magnetic field $B$. The axion-photon coupling of strength $g_{A\gamma\gamma}=(0.68\pm 0.02)\times 10^{-10}~{}{\rm GeV^{-1}}$ [12] changes the photon’s Green’s function [13, 14]
$\Delta D_{+}(x,x^{\prime})=m_{A}^{2}c^{4}g_{A\gamma\gamma}\tilde{\phi}(q_{A})D_{+}(x,x^{\prime}),$ (3)
and this change in the photon propagator produces the shift of the spin precession frequency [15]:
$\frac{\Delta a_{\mu}}{a_{\mu}}$ $=$ $m_{A}^{2}c^{4}g_{A\gamma\gamma}\tilde{\phi}_{A}(q_{A})$ (4)
$=$ ±$\pm m_{A}c^{2}g_{A\gamma\gamma}\sqrt{\rho_{A}N_{\mu}V_{\mu}T_{\mu}/\hbar}$
where, the axion field strength $|\tilde{\phi}_{A}(q_{A})|^{2}\approx\phi_{A}^{2}N_{\mu}V_{\mu}T_{\mu}/\hbar$ is fixed by the axion energy density $\rho_{A}=\phi_{A}^{2}m_{A}^{2}c^{4}$, the volume $V_{\mu}$ of the muon beam, the lifetime $T_{\mu}$ of the muon in the axion rest-frame, and the average number of muons $N_{\mu}$ during spin precession. The best astrophysical estimate for the energy density of dark matter in the local halo is $\rho_{A}=(0.30\pm 0.03)~{}{\rm GeV/cm^{3}}$ [16].
The sign of the axion-driven shift $\Delta a_{\mu}$ of the muon spin precession frequency switches along with the axion field $\tilde{\phi}_{A}(q_{A})$ under spatial inversion $P$, under time-reversal $T$, and under the combinations $CP$ and $CT$ with charge conjugation $C$ [17]. The size of the shift scales with muon beam parameters [18]
$\Delta a_{\mu}$ $=$ $(171\pm 10)\times 10^{-11}\left(\frac{N_{\mu}}{1930}\right)^{1/2}\left(\frac{\sigma_{x}}{1.78~{}{\rm cm}}\right)^{1/2}$
$\times$ $\left(\frac{\sigma_{y}}{1.25~{}{\rm cm}}\right)^{1/2}\left(\frac{R}{711~{}{\rm cm}}\right)^{1/2}\left(\frac{T_{\mu}}{64.4~{}{\rm\mu sec}}\right)^{1/2}$
where, $\sigma_{x}$ is the width of the muon beam, $\sigma_{y}$ is its height, and $R$ is the radius of the storage ring. Combining the shifts found from Eq. (Weighing the Axion with Muon Haloscopy) using the muon beam parameters for the two experiments (FNAL E989 [2, 3] and BNL E821[4]) that dominate the experimental world average $a_{\mu}({\rm Exp})$, we find the weighted average shift
$\Delta a_{\mu}({\rm Exp})=(171\pm 10)\times 10^{-11},$ (6)
where, we have used the same weights for $\Delta a_{\mu}({\rm Exp})$ as the weights entering $a_{\mu}({\rm Exp})$ [18]. This shift roughly resolves the $3.4~{}\sigma$ tension between $a_{\mu}({\rm Exp})$ and the standard model value $a_{\mu}({\rm SM})$[19, 20, 21, 22, 23, 24]:
$a_{\mu}({\rm Exp})-a_{\mu}({\rm SM})=(140\pm 41)\times 10^{-11},$ (7)
where, we have used the recently emerged consensus value $a_{\mu}^{\rm HLO}({\rm SM})=(704.1\pm 3.0)\times 10^{-10}$ [25] for the leading order hadronic contribution to the muon spin precession frequency within the Standard Model [26].
A remarkable consequence of this resolution of the tension is that it can be tested in a relatively straightforward fashion. Flipping the magnetic storage ring over the muon beamline, such that the muons circulate with the opposite sense around the ring, would yield a spin precession frequency in tension with the present experimental world average by $2\Delta a_{\mu}({\rm Exp})/a_{\mu}({\rm Exp})=2\times(2.71\pm 0.17)~{}{\rm BNL}=(5.4\pm 0.3)~{}{\rm BNL}$, where $1~{}{\rm BNL}=0.54~{}{\rm ppm}$ [26] is the precision achieved at BNL E821 [4] and is the precision that can be achieved at FNAL E989 [2, 3] in a data collecting run with the flipped ring lasting roughly one month. The uncertainty of the present world average, roughly $0.4~{}{\rm BNL}$ [2, 26], combines with the statistics-dominated uncertainty of such a short run to give an expected significance of the tension $(5.1\pm 0.3)\sigma$ that surpasses the conventional threshold for the discovery of a new phenomenon in particle physics, where $\sigma\approx 1~{}{\rm BNL}$ is the uncertainty of the comparison.
Making Axions in Polar Materials with Infrared— Axions from the local dark matter halo of the galaxy can be seen in the mid-infrared electromagnetic response of a wide range of polar materials such as silicon nitride ${\rm Si_{3}N_{4}}$ [27]. The incoming mid-infrared light with frequency $\omega_{A\theta}=\omega_{A}-\omega_{\theta}$ makes axions within the polar material with energy $\hbar\omega_{A}=m_{A}c^{2}=h\times(122.7\pm 1.0)~{}{\rm THz}$ by absorbing a transverse optical phonon at the far-infrared frequency $\omega_{\theta}\approx 2.4~{}{\rm THz}$ [27]. The axions leave the material and create a dip in the infrared transmission through the sample at the axion resonance frequency $\omega_{A\theta}$.
Measuring the mid-infrared frequency $\omega_{A\theta}=\omega_{A}-\omega_{\theta}$ at which the polar material makes axions and the far-infared frequency $\omega_{\theta}$ of the transverse optical phonons, one finds the rest-mass energy of the axion $m_{A}c^{2}=\hbar\omega_{A}=\hbar(\omega_{A\theta}+\omega_{\theta}).$ The size of the dip in mid-infrared transmission through a sample of the polar material on the axion resonance at $\omega_{A\theta}$ is set by the infrared absorption coefficient [28]:
$\alpha(\omega_{A\theta})$ $=$ $\alpha(\omega_{\theta})\frac{\Gamma_{A\theta}}{\Gamma_{\theta}}\frac{\omega_{\theta}}{\omega_{A\theta}},$ (8)
where $\alpha(\omega_{\theta})$ is the absorption coefficient of the phonon, $\Gamma_{A\theta}\approx 2\pi c/t$ gives the width of the axion resonance in terms of the thickness of the sample $t$, and $\Gamma_{\theta}$ is the width of the phonon resonance.
Using typical parameters for a polar material such as silicon nitride ${\rm Si_{3}N_{4}}$, we find the mid-infrared absorption coefficient at $\omega_{A\theta}$ from the axion resonance [27]
$\alpha(\omega_{A\theta})$ $=$ $4.0\times\alpha(\omega_{\theta})\left(\frac{t}{5~{}{\rm\mu m}}\right)^{-1}\left(\frac{\Gamma_{\theta}/2\pi c}{10~{}{\rm cm^{-1}}}\right)^{-1}$ (9)
$\times$ $\left(\frac{\omega_{\theta}/2\pi c}{80~{}{\rm cm^{-1}}}\right)\left(\frac{\omega_{A\theta}/2\pi c}{4016~{}{\rm cm^{-1}}}\right)^{-1}.$
The mid-infrared absorption with coefficient $\alpha(\omega_{A\theta})\approx 4\alpha(\omega_{\theta})$ in polar materials such as silicon nitride ${\rm Si_{3}N_{4}}$ at frequency $\omega_{A\theta}=\omega_{A}-\omega_{\theta}$ due to creation of axions can thus be seen with readily available samples and infrared spectrometers provided the spectrometer sees the dip with absorption coefficient $\alpha(\omega_{\theta})$ in the far-infrared transmission spectrum at frequency $\omega_{\theta}$ from the transverse optical phonon in the sample [27].
Figure 2: Tentative observation of the axion resonance in the Raman light scattering spectrum of synthetic quartz crystal (${\rm SiO_{2}}$, RRUFF deposition R110108) [30, 29] at room temperature and ambient pressure. The axion rest-mass (See Eq. 2) gives the window (vertical lines) for the frequency shift $\omega_{A}=m_{A}c^{2}/\hbar=2\pi c\times(4096\pm 32)~{}{\rm cm^{-1}}$ of visible light that makes axions within the polar material. A broad band of scattered light intensity is observed with the peak frequency shift falling within the predicted axion resonance window.
Axion Resonance in Quartz Raman Spectroscopy— A tentative observation of the axion resonance in the mid-infrared electrodynamics response of quartz (${\rm SiO_{2}}$) [29] has been reported by the RRUFF project [30] using Raman light scattering spectroscopy (Fig. 2). The scattered light intensity has a broad peak centered on the frequency shift $\omega_{A}({\rm SiO_{2}})=2\pi c\times(4080\pm 4)~{}{\rm cm^{-1}}$ which falls within the predicted window $\omega_{A}=m_{A}c^{2}/\hbar=2\pi c\times(4096\pm 32)~{}{\rm cm^{-1}}$. Quartz has no known Raman-active vibration modes in this frequency range and chemical analysis suggests that the sample has insufficient concentration of luminescent impurities to account for the strength of the observed resonance [29].
Assuming that visible light incident on the quartz crystal makes axions within the sample, the frequency shift of the scattered light where the intensity peaks $\omega_{A}({\rm SiO_{2}})$ can be interpreted in terms of the rest-mass energy of the axion:
$m_{A}({\rm SiO_{2}})c^{2}=\hbar\omega_{A}({\rm SiO_{2}})=(0.5059\pm 0.0005)~{}{\rm eV}.$ (10)
This agrees with the estimate in Eq. (2) but is roughly ten times more precise. If this interpretation is confirmed, then it would provide strong support for the novel physical picture of dark matter in terms of axions with number density $n_{A}=6n_{\gamma}$.
Indeed, the rest-mass energy of the axion from quartz in Eq. (10), when combined with the axion number density $n_{A}=6n_{\gamma}=(2464.2\pm 1.6)~{}{\rm cm^{-3}}$ [8] suggested by the novel physical picture of dark matter, gives the cosmological energy density parameter for axions:
$\Omega_{A}h^{2}({\rm SiO_{2}})=\frac{8\pi G}{3}\frac{h^{2}}{H_{0}^{2}}n_{A}m_{A}=(0.11831\pm 0.00008),$ (11)
where, $H_{0}/h=100~{}{\rm km/(sec\cdot Mpc)}=1/(9.778~{}{\rm Gyr})$ [8] gives the conventional parameterization of the present expansion rate of the universe. This agrees with the best estimate for the cosmological energy density parameter of dark matter $\Omega_{A}h^{2}=(0.11882\pm 0.00086)$ [10]. However, it is ten times times more precise.
Conclusion— We conclude that the tensions between measurements of muon spin precession and standard model calculations are due to axions from the local dark matter halo of the galaxy. The resolution of the muon tension rests on a novel physical picture for the nature of dark matter. The picture leads to percent level predictions for the mass of the axion and for its coupling to electromagnetic fields which can be tested using a tabletop experiment with readily available materials and devices.
Acknowledgements— The author acknowledges the support of Daria Mazura.