### v38

# Weighing the Axion with Muon Haloscopy

Abstract
Recent measurements of muon spin precession confirm a long-standing tension with the standard model of particle physics. We argue that axions form the dark matter in the universe and that axions from the local dark matter halo of the galaxy are responsible for the tension. The argument gives a percent level prediction for the mass of the axion that can be tested by measuring the mid-infrared electrodynamic response of a wide range of polar materials. A broad resonance in the publicly available Raman spectrum from quartz is consistent with the creation of axions with the predicted mass. The difference in the frequency of a pair of absorption lines in the published infrared spectrum of polystyrene gives an estimate of the mass of the axion to 44 parts per million precision that is consistent with the predicted value.
Introduction—The spin of the muon precesses in a magnetic field $B$ at frequency $\omega_{a}=a_{\mu}(e/m_{\mu})B$, where, $a_{\mu}\approx\alpha/(2\pi)\approx 1.16\times 10^{-3}$ is the electromagnetic leading order contribution to spin precession, and $e/m_{\mu}=2\pi\times 135~{}{\rm MHz/T}$ is the charge to mass ratio of the muon [1, 2, 3, 4, 5]. In this Letter, we show that axions in the local dark matter halo of the galaxy shift $a_{\mu}$ from the standard model value. The demonstration rests on a novel physical picture for the nature of dark matter which leads to a percent level prediction for the mass of the axion.
Nature of Dark Matter—We start by expressing the axion $A_{M}$ in terms of quarks and leptons $M^{P}_{H}$ with helicity $H=L,R$ and chirality check $P=+,-$ [6]:
$A_{M}$ $=$ $M_{L}^{-}\overline{M}_{R}^{-}-M_{R}^{+}\overline{M}_{L}^{+}-M_{L}^{+}\overline{M}_{R}^{+}+M_{R}^{-}\overline{M}_{L}^{-}$ (1)
Here, $M$ runs over the twelve known “flavors” of quarks and leptons while the chirality check $P$ is $+$ for quarks and leptons that do not feel the weak nuclear force and -- for those that do. In the standard model, we find $M_{L}^{-}$ and $M_{R}^{+}$, but neither $M_{L}^{+}$ nor $M_{R}^{-}$, where, $L$ is for left-handed and $R$ is for right-handed helicity [7].
Next, we suppose that axions $A_{M}$ of each kind and photons $\gamma_{H}$ of each helicity formed in equal numbers in the early universe prior to baryogenesis. This fixes the ratio $n_{A}/n_{\gamma}=|M|/|H|=6$ of axions to photons and gives the axion rest-mass energy [8]:
$m_{A}c^{2}=2.70~{}kT_{\gamma}\left(\frac{n_{A}}{n_{\gamma}}\right)^{-1}\frac{\Omega_{A}h^{2}}{\Omega_{\gamma}h^{2}}=(0.508\pm 0.004)~{}{\rm eV},$ (2)
where, $T_{\gamma}=(2.7255\pm 0.0006)~{}{\rm K}$ is the present photon temperature [9], $\Omega_{\gamma}h^{2}=(2.473\pm 0.002)\times 10^{-5}$ is the photon energy density parameter and $\Omega_{A}h^{2}=0.11882\pm 0.00086$ is that of dark matter [10].
Muon Haloscopy—Axions in the local dark matter halo shift muon spin precession [3, 4, 5]. Acting as a background field $\tilde{\phi}_{A}(q_{A})$ with on-shell four-momentum $q_{A}^{2}=(m_{A}c)^{2}\ll(m_{\mu}c)^{2}$, the axion $A$ couples to the virtual photon dressing the vertex of the positive muon $\mu^{+}$ and the magnetic field $B$. The axion-photon coupling of strength $g_{A\gamma\gamma}=(0.68\pm 0.02)\times 10^{-10}~{}{\rm GeV^{-1}}$ [12] changes the photon’s Green’s function [13, 14]
$\Delta D_{+}(x,x^{\prime})=m_{A}^{2}c^{4}g_{A\gamma\gamma}\tilde{\phi}(q_{A})D_{+}(x,x^{\prime}),$ (3)
and this change in the photon propagator produces the shift of the spin precession frequency [15]:
$\frac{\Delta a_{\mu}}{a_{\mu}}$ $=$ $m_{A}^{2}c^{4}g_{A\gamma\gamma}\tilde{\phi}_{A}(q_{A})$ (4)
$=$ ±$\pm m_{A}c^{2}g_{A\gamma\gamma}\sqrt{\rho_{A}N_{\mu}V_{\mu}T_{\mu}/\hbar}$
where, the axion field strength $|\tilde{\phi}_{A}(q_{A})|^{2}\approx\phi_{A}^{2}N_{\mu}V_{\mu}T_{\mu}/\hbar$ is fixed by the axion energy density $\rho_{A}=\phi_{A}^{2}m_{A}^{2}c^{4}$, the volume $V_{\mu}$ of the muon beam, the lifetime $T_{\mu}$ of the muon in the axion rest-frame, and the average number of muons $N_{\mu}$ during spin precession. The best astrophysical estimate for the energy density of dark matter in the local halo is $\rho_{A}=(0.30\pm 0.03)~{}{\rm GeV/cm^{3}}$ [16].
The sign of the axion-driven shift $\Delta a_{\mu}$ of the muon spin precession frequency switches along with the axion field $\tilde{\phi}_{A}(q_{A})$ under spatial inversion $P$, under time-reversal $T$, and under the combinations $CP$ and $CT$ with charge conjugation $C$ [17]. The size of the shift scales with muon beam parameters [18]
$\Delta a_{\mu}$ $=$ $(171\pm 10)\times 10^{-11}\left(\frac{N_{\mu}}{1930}\right)^{1/2}\left(\frac{\sigma_{x}}{1.78~{}{\rm cm}}\right)^{1/2}$
$\times$ $\left(\frac{\sigma_{y}}{1.25~{}{\rm cm}}\right)^{1/2}\left(\frac{R}{711~{}{\rm cm}}\right)^{1/2}\left(\frac{T_{\mu}}{64.4~{}{\rm\mu sec}}\right)^{1/2}$
where, $\sigma_{x}$ is the width of the muon beam, $\sigma_{y}$ is its height, and $R$ is the radius of the storage ring. Using the muon beam parameters for the two experiments (FNAL E989 [2, 3] and BNL E821[4]) that dominate the experimental world average $a_{\mu}({\rm Exp})$, we find the weighted average shift
$\Delta a_{\mu}({\rm Exp})=(171\pm 10)\times 10^{-11},$ (6)
where, we have used the same weights for $\Delta a_{\mu}({\rm Exp})$ as the weights entering $a_{\mu}({\rm Exp})$ [18]. This shift roughly resolves the $3.4~{}\sigma$ tension between $a_{\mu}({\rm Exp})$ and the standard model value $a_{\mu}({\rm SM})$[19, 20, 21, 22, 23, 24]:
$a_{\mu}({\rm Exp})-a_{\mu}({\rm SM})=(140\pm 41)\times 10^{-11},$ (7)
where, we have used the recently emerged consensus value $a_{\mu}^{\rm HLO}({\rm SM})=(704.1\pm 3.0)\times 10^{-10}$ [25] for the leading order hadronic contribution to the muon spin precession frequency within the Standard Model [26].
Flipping the magnetic storage ring over the muon beamline, such that the muons circulate with the opposite sense around the ring, would yield a spin precession frequency in tension with the present experimental world average by $2\Delta a_{\mu}({\rm Exp})/a_{\mu}({\rm Exp})=2\times(2.71\pm 0.17)~{}{\rm BNL}=(5.4\pm 0.3)~{}{\rm BNL}$. Here $1~{}{\rm BNL}=0.54~{}{\rm ppm}$ [26] is the precision achieved at BNL E821 [4] and is the precision that can be achieved at FNAL E989 [2, 3] in a data collecting run with the flipped ring lasting roughly one month. The uncertainty of the present world average, roughly $0.4~{}{\rm BNL}$ [2, 26], combines with the statistics-dominated uncertainty of such a short run to give an expected $(5.1\pm 0.3)\sigma$ significance of the tension.
Figure 1: Tentative observation of the axion resonance in the Raman light scattering spectrum of synthetic quartz crystal (${\rm SiO_{2}}$, RRUFF deposition R110108) [28, 27] at room temperature and ambient pressure. The axion rest-mass (See Eq. 2) gives the window (vertical lines) for the frequency shift $\omega_{A}=m_{A}c^{2}/\hbar=2\pi c\times(4096\pm 30)~{}{\rm cm^{-1}}$ of visible light that makes axions within the polar material. A broad band of scattered light intensity is observed with the peak frequency shift falling within the predicted axion resonance window.
Axion Resonance in Quartz Raman Spectroscopy— A tentative observation of the axion resonance in the mid-infrared electrodynamics response of quartz (${\rm SiO_{2}}$) [27] has been reported by the RRUFF project [28] using Raman light scattering spectroscopy (Fig. 1). The scattered light intensity has a broad peak centered on the frequency shift $\omega_{A}({\rm SiO_{2}})=2\pi c\times(4080\pm 4)~{}{\rm cm^{-1}}$ which falls within the predicted window $\omega_{A}=m_{A}c^{2}/\hbar=2\pi c\times(4096\pm 30)~{}{\rm cm^{-1}}$. Quartz has no known Raman-active vibration modes in this frequency range and chemical analysis suggests that the sample has insufficient concentration of luminescent impurities to account for the strength of the observed resonance [27].
Assuming that visible light incident on the quartz crystal makes axions within the sample using the dot product ${\bf E}_{\rm IN}\cdot{\bf B}_{\rm OUT}$ [12] of the electric field ${\bf E}_{\rm IN}$ of the incoming photon and the magnetic field ${\bf B}_{\rm OUT}$ of the outgoing photon, then the frequency shift of the scattered light where the intensity peaks $\omega_{A}({\rm SiO_{2}})$ gives the rest-mass energy of the axion:
$m_{A}({\rm SiO_{2}})c^{2}=\hbar\omega_{A}({\rm SiO_{2}})=(0.5059\pm 0.0005)~{}{\rm eV}.$ (8)
This agrees with the estimate in Eq. (2) but is about 8 times more precise. If this interpretation is confirmed, then it would provide strong support for the novel physical picture of dark matter in terms of axions with number density $n_{A}=6n_{\gamma}$.
Indeed, the rest-mass energy of the axion from quartz in Eq. (8), when combined with the axion number density $n_{A}=6n_{\gamma}=(2464.2\pm 1.6)~{}{\rm cm^{-3}}$ [8] suggested by the novel physical picture of dark matter, gives the cosmological energy density parameter for axions:
$\Omega_{A}h^{2}({\rm SiO_{2}})=\frac{8\pi G}{3}\frac{h^{2}}{H_{0}^{2}}n_{A}m_{A}=(0.11831\pm 0.00008),$ (9)
where, $H_{0}/h=100~{}{\rm km/(sec\cdot Mpc)}=1/(9.778~{}{\rm Gyr})$ [8] gives the conventional parameterization of the present expansion rate of the universe. This agrees with the best estimate for the cosmological energy density parameter of dark matter $\Omega_{A}h^{2}=(0.11882\pm 0.00086)$ [10]. However, it is around 10 times more precise.
Axion Mass from Infrared Spectroscopy— An even more precise determination of the mass of the axion to 44 parts per million follows from the analysis of the published near-infrared transmittance spectrum of the material polystyrene taken from clear plastic cases used to store compact disks (CD) [29]. An incoming near-infrared photon at frequency $\nu_{A\theta}$ makes an axion at the Compton frequency $\nu_{A}({\rm CD})=m_{A}({\rm CD})/(hc)$ while exciting the infrared-active phonon mode at frequency $\nu_{\theta}$ within the polystyrene. The source of the axion is the dot product ${\bf E}_{\rm NIR}\cdot{\bf B}_{\rm FIR}$ of the near-infrared electric field ${\bf E}_{\rm NIR}$ from the incoming photon and the far-infrared magnetic field ${\bf B}_{\rm FIR}$ from the phonon [12].
Remarkably, a pair of absorption lines in the near-infrared transmittance spectrum of polystyrene from CD cases [29] excite the same infrared-active phonon with frequency $\nu_{\theta}$, yet they do so by making a pair of axions or just one axion [30]:
$2\nu_{A}({\rm CD})+\nu_{\theta}$ $=$ $(8742.84\pm 0.16\pm 0.055)~{}{\rm cm^{-1}}$ (10)
$\nu_{A}({\rm CD})+\nu_{\theta}$ $=$ $(4663.27\pm 0.013\pm 0.055)~{}{\rm cm^{-1}}$
where the first uncertainty is statistical and the second is systematic. Taking the difference of this pair of frequencies where the transmitted near-infrared light intensity dips when passing through the polystyrene from CD cases, we can weigh the axion with infrared spectroscopy to 44 ppm precision [31]
$\nu_{A}({\rm CD})$ $=$ $(4079.57\pm 0.18)~{}{\rm cm^{-1}}~{}[44~{}{\rm ppm}].$ (11)
The resulting mass of the axion $m_{A}c^{2}({\rm CD})=hc\times\nu_{A}=(0.50580\pm 0.00002)~{}{\rm eV}$ agrees with the mass $m_{A}({\rm SiO_{2}})$ from the Raman spectroscopy of quartz in Eq. (8), but it is roughly 20 times more precise.
Conclusion— We conclude that the tensions between measurements of muon spin precession and standard model calculations are due to axions from the local dark matter halo of the galaxy. The resolution of the muon tension rests on a novel physical picture for the nature of dark matter. The picture leads to percent level predictions for the mass of the axion and for its coupling to electromagnetic fields which can be tested using tabletop experiments with readily available materials and devices, including the Raman spectroscopy of a synthetic quartz crystal and the infrared spectroscopy of polystyrene from clear plastic cases for compact discs.
Acknowledgements— The author acknowledges the support of Daria Mazura.